a man and a women appear in an interview for two vacancies in the same post. the probability of mans selection is 1/4 and that of women s selectionis 1/3 the probability that none of them will be selected is- (a) 1/2 (b) 1/12 (c) 1/4 (d) none of these
1
Expert's answer
2012-05-31T08:10:45-0400
P{none of them will be selected} = P{man will not be selected}·P{woman will not be selected} = (1-1/4)(1-1/3) = 3/4·2/3 = 2/4 = 1/2.

So, the right answer is (a).